The U.S. Department of Health and Human Services reports that there were 250,000

babies born to teenagers aged 15-18 in 2004. Of these 250,000 births, 7% were to

15-year-olds, 17% were to 16-year-olds, 29% were to 17-year-olds, and 47% were to

18-year-olds. LetX be the age of a randomly selected mother between the ages of 15

to 18. Find the mean, variance, and standard deviation of X
Answer #1

Mean = E(X) = 0.07 * 15 + 0.17 * 16 + 0.29 * 17 + 0.47 * 18 = 17.16

E(X2) = 0.07 * 152 + 0.17 * 162 + 0.29 * 172 + 0.47 * 182 = 295.36

Variance = E(X2) - (E(X))2 = 295.36 - 17.162 = 0.8944

Standard deviation = sqrt(0.8944) = 0.9457